OpenStudy (anonymous):

Evaluate (medal given for steps shown ^_^) \[ \huge \int_{0}^{1/4} \frac{sin^{-1}(2x)}{\sqrt{1-4x^2}} dx \]

OpenStudy (anonymous):

Okay, so we need to do a u-substitution right? So let: \[u=\arcsin(2x) \implies du = \frac{2 dx}{\sqrt{1-4x^2}} \implies \frac{du}{2}=\frac{dx}{\sqrt{1-4x^2}}\] Therefore the integral becomes: \[\int\limits_{u_1}^{u_2} \frac{u}{2}du\] Evaluate your bounds: arcsin(2(1/4))=arcsin(1/2)=pi/6 (assuming typical branch cut on arcsin from -pi/2 to pi/2) and arcsin(2(0))=0 So we have: \[\int\limits_0^{\pi/6}\frac{u}{2}du=\left[ \frac{u^2}{4}\right]_0^{\pi/6}=\frac{\pi^2}{144}\]

OpenStudy (anonymous):

Excellent, ty